OpenStudy (anonymous):

Could someone please solve and explain the following problem? 5x^2 + 1 = 37 Thanks!

OpenStudy (anonymous):

plus minus 6 divided by the square root of 5.

OpenStudy (anonymous):

Use Quadratic formula, you will get \( \large \left\{\left\{x= -\frac{6}{\sqrt{5}}\right\},\left\{x = \frac{6}{\sqrt{5}}\right\}\right\} \)

OpenStudy (anonymous):

Here are the steps.... 5x^2 + 1 - 37 = 0 implies 5x^2 - 36 = 0 implies 5x^2 = 36 implies x^2 = 36/5 implies x = plus minus 6/root 5

OpenStudy (anonymous):

Ok, that was my answer too. The textbook said it was 6 times the root of 5 over 5...

OpenStudy (anonymous):

W\[6\sqrt{5}/5 * \sqrt{5}/\sqrt{5} = 6*5/5\sqrt{5} = 6/\sqrt{5}\]

OpenStudy (anonymous):

^^ ohhhhh ok. i'm just glad i know i wasn't missing some really big step. xD thanks for you help!

OpenStudy (anonymous):

Welcome